OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

m=(1-2)/(4-5)=1, y-2=-1(x-5), y-2=-x+5, x+y=7